OpenStudy (anonymous):

PLEASE HELP The rotating light on a lighthouse is 400 feet from a cliff. It completes one rotation every 10 seconds. The equation representing the distance, d, in feet that the center of the circle of light is from the lighthouse is d(t) = 400sec (πt/5). What is the period of d(t)? (Enter only the number.)

OpenStudy (anonymous):

Turing test do you know?

OpenStudy (turingtest):

If your just asking about the period I don't see why it gives the whole thing about distance. \[\sec x={1\over \cos x}\]and these functions in the form of \[A \cos(Bx+C)\] where the period is\[ {2\pi \over \left| B \right|}\]so the period of your function is \[{2\pi \over{\pi \over 5}}=10\]I just don't see what to do with the rest of the info.

OpenStudy (anonymous):

ok thanks IT WAS RIGHT

OpenStudy (turingtest):

But why all the extra info... seems weird :/